### v1

@@ 2.4 The relationships between linear consensus algorithm and effective resistance
The proof is analogous to [19].
Let $\Phi(P):=\Phi_{n}(P)=n\varPi P$, where $n$ denotes the number of rows of $P$. Then, we can see a clear relationship between the effective resistance and the random walk with the transition matrix $P$ [9].
Lemma 2.5.
Let $\mathcal{G}$ be a resistor network with the conductance matrix $C:=\Phi(P)$. Consider the Markov chain with the transition matrix $P$, and let $p_{\mathrm{esc}}$ be the probability that the walker on the Markov chain starting at $a$ returns $a$ before reaching $b$. Then $p_{\mathrm{esc}}=\frac{1}{\mathcal{R}_{ab}\sum_{x}{C_{ax}}}$.
There are also random walk interpretations of the electrical network other than resistance, such as the voltage and the current of the network (see [9]). In the analogy of random walk, a matrix called Green matrix [19] plays an important role. The Green matrix of consensus matrix $P$ is defined as:
@@
This matrix is also called the fundamental matrix [1]. For a consensus matrix $P$, this matrix is well defined (see Chapter 11.4 in [14]).
@@
Because $P^{t}-\bm{1}\bm{\pi}=(P-\bm{1}\bm{\pi})^{t}$ for $t\geq 1$, (7) can be reformulated as follows:
(8) $G(P)+\bm{1}\bm{\pi}=I+\sum_{t\geq 1}(P-\bm{1}\bm{\pi}^{\mathsf{T}})^{t}.$
@@
(9) $G(P)+\bm{1}\bm{\pi}=(I-P+\bm{1}\bm{\pi}^{\mathsf{T}})^{-1}.$
A fundamental matrix often refers not to $G(P)$, but to the right hand of (8) or (9) (e.g. [14, 18]).
Now we show another representation of effective resistance using the Green matrix $G(P)$.
Lemma 2.7.
For a reversible consensus matrix $P$, let $C=\Phi(P)$. Then,
$\mathcal{R}_{ab}(C)=\frac{1}{n}(\bm{e}_{a}-\bm{e}_{b})^{\mathsf{T}}G(P)\varPi^{-1}(\bm{e}_{a}-\bm{e}_{b}).$
The proof can be found in [19]. If we start the discussion from a conductance matrix $C$, we can obtain the following corollary.
Corollary 2.8.
For a conductance matrix $C$, let $P=\Psi(C)$. Then,
$\mathcal{R}_{ab}(C)=(\bm{e}_{a}-\bm{e}_{b})^{\mathsf{T}}G(P)\operatorname{diag}(C\bm{1})^{-1}(\bm{e}_{a}-\bm{e}_{b}).$
The proparties in this section will be used to relate $J(P)$ and the effective resistance of the graph. In particular, $G(P)$ plays an important role in evaluating $J(P)$.
@@ 3 Performance analysis of linear consensus algorithm using effective resistance
In this section, we give upper and lower bounds for $J(P)$ using the concept of effective resistance. A bound for $J(P)$ has been known for reversible $P$ [19, 18], but not for nonreversible cases. To provide bounds for nonreversible $P$, we use a reversible matrix $P^{*}P$, where $P^{*}:=\varPi^{-1}P^{\mathsf{T}}\varPi$. Our result is a generalization of [19], which relates $J(P)$ and ¯$\bar{\mathcal{R}}(\Phi(P^{2}))$.
@@ 3.1 Properties of $P^{*}P$
Because $P$ is normal, $(I-P)(I-P^{\mathsf{T}})=(I-P^{\mathsf{T}})(I-P)$. Multiplying $\bm{1}$ from the right, we obtain $(I-P)(I-P^{\mathsf{T}})\bm{1}=\bm{0}$, which means $(I-P^{\mathsf{T}})\bm{1}\in\ker(I-P)$. Therefore, there exists some real value $a$ that satisfies $(I-P^{\mathsf{T}})\bm{1}=a\bm{1}$, because $\ker(I-P)=\operatorname{span}\{\bm{1}\}$. Then we multiply $\bm{1}^{\mathsf{T}}$ from the left and obtain $\bm{0}=an$. This means that $a=0$ and $(I-P^{\mathsf{T}})\bm{1}=\bm{0}$. Thus, $P$ is doubly-stochastic and the invariant measure $\bm{\pi}$ is $\frac{1}{n}\bm{1}$. Therefore, we obtain $P^{*}=P$ and $P^{*}P=P^{\mathsf{T}}P=PP^{\mathsf{T}}=PP^{*}$.
@@
Inclusion properties among classes of consensus matrices are shown in Figure 1. Notice that there are also a matrix $P$ which is not reversible, nor normal, but satisfies $P^{*}P=PP^{*}$, such as the following example (11). Thus, the condition $P^{*}P=PP^{*}$ is a weeker assumption than reversibility or normality.
@@
3.2 A bound for $J(P)$ and $J_{\mathrm{w}}(P)$ using ¯$\bar{\mathcal{R}}(C)$
@@
(12) ¯$\bar{\mathcal{R}}_{\mathrm{w}}(C):=\frac{1}{2}\bm{\pi}^{\mathsf{T}}\mathcal{R}(C)\bm{\pi}=\frac{1}{2}\sum_{(u,v)\in V\times V}\mathcal{R}_{uv}(C)\pi_{u}\pi_{v},$
@@
¯$\bar{\mathcal{R}}_{\mathrm{w}}(C)\leq\frac{1}{2}\pi_{\max}^{2}\sum_{(u,v)\in V\times V}\mathcal{R}_{uv}(C)=n^{2}\pi_{\max}^{2}\bar{\mathcal{R}}(C),$
@@
¯$\bar{\mathcal{R}}_{\mathrm{w}}(C)\geq\frac{1}{2}\pi_{\min}^{2}\sum_{(u,v)\in V\times V}\mathcal{R}_{uv}(C)=n^{2}\pi_{\min}^{2}\bar{\mathcal{R}}(C).$
@@
¯$\bar{\mathcal{R}}_{\mathrm{w}}(C)=\frac{1}{n}\operatorname{tr}G(P),$
@@
Using ¯$\bar{\mathcal{R}}_{\mathrm{w}}(C)$, we prove Theorem 3.5.
@@
because $J(P)=\operatorname{tr}G_{\mathrm{w}}$ by definition. Since
@@
Lemmas 3.6 and 3.8 implies that
@@
Remark 3.11.
It seems to be hard to give a lower bound for $\operatorname{tr}\left((P^{*})^{t}P^{t}\right)$ using $\operatorname{tr}(P^{*}P)^{t}$, because $\operatorname{tr}(P^{*}P)^{t}$ can be much larger than $\operatorname{tr}(P^{*}P)^{t}$. Let us consider the following consensus matrix:
$P^{t}-\bm{1}\bm{\pi}$ converges exponentially with respect to the second largest eigenvalue of $P$. The eigenvalues of $P_{\varepsilon}$ are $1$ and $\alpha_{\pm}:=-\left(\frac{1}{4}-\varepsilon\right)\pm\frac{\mathrm{i}}{2}\sqrt{\frac{7}{4}-2\varepsilon}$. The eigenvalues of $P_{\varepsilon}^{*}$ are the same due to similarity. Therefore, $(P_{\varepsilon}^{*})^{t}P_{\varepsilon}^{t}-\bm{1}\bm{\pi}$ converges exponentially to some real matrix with respect to the rate near $1/2$.
On the other hand, $P^{*}P$ has the eigenvalue $1-O(\varepsilon)$. This means that if $\varepsilon$ is small enough, the convergence of $(P^{*}P)^{t}$ is very slow, and its trace can be far larger than $\operatorname{tr}(P_{\varepsilon}^{*})^{t}P_{\varepsilon}^{t}$.
@@ 3.3 A bound for $J(P)$ using the graph structure of network
In this section, we establish bounds for $J(P)$ using the structure of undirected associated graph $\mathcal{G}(P)$, the maximum and minimum elements of $\bm{\pi}$, and the minimum element of $P$. Notice that we do not consider the edge weights of $\mathcal{G}(P)$ and focus only on its graph topology of $\mathcal{G}(P)$.
@@
Let $P$ be a consensus matrix with invariant measure $\pi$, and let $\mathcal{G}(P)$ be the associated undirected graph with $P$. Then it holds that
@@
where $\pi_{\min}$ and $\pi_{\max}$ are respectively the minimum and maximum entries of $\bm{\pi}$, $p_{\min}$ is the minimum nonzero entry of $P$, and ¯$\bar{\mathcal{R}}(\mathcal{G}(P))$ is the average effective resistance of a network, associated with $\mathcal{G}(P)$, such that each edge has a unit conductance. Moreover, if $PP^{*}=P^{*}P$, then
$J(P)\geq\frac{\pi_{\min}^{3}n}{4p_{\max}^{2}\delta_{\mathrm{in}}^{2}\pi_{\max}^{2}}\bar{\mathcal{R}}(\mathcal{G}(P)),\quad J_{\mathrm{w}}(P)\geq\frac{\pi_{\min}^{3}n^{2}}{4p_{\max}^{2}\delta_{\mathrm{in}}^{2}\pi_{\max}}\bar{\mathcal{R}}(\mathcal{G}(P)).$
where $p_{\max}$ is the maximum nonzero entry of $P$, and $\delta_{\mathrm{in}}$ is the maximum indegree (excluding self loops) of the associated directed graph $\mathcal{G}_{\mathrm{dir}}(P)$.
By Theorem 3.5, it is enough for the proof to show the following lemma, which is the generalization of Lemma 5.9 in [19].
@@
¯$\bar{\mathcal{R}}(C_{P^{*}P})\leq\frac{1}{n\pi_{\min}p_{\min}^{2}}\bar{\mathcal{R}}(\mathcal{G}(P)).$
@@
¯$\bar{\mathcal{R}}(C_{P^{*}P})\geq\frac{1}{4n\pi_{\max}\delta_{\mathrm{in}}p_{\max}^{2}}\bar{\mathcal{R}}(\mathcal{G}(P)).$
Since the associated graph of the consensus matrix $P$ can be directed, we evaluate ¯$\bar{\mathcal{R}}(C_{P^{*}P})$ using the maximum indegree $\delta_{\mathrm{in}}$ of the graph nodes of $\mathcal{G}_{\mathrm{dir}}(P)$, unlike previous research [19].
We employ the undirected graph $\mathcal{G}(P)$ instead of the directed graph $\mathcal{G}_{\mathrm{dir}}$, because the following property is essential for proving Theorem 3.12.
@@
Using Lemma 3.14, we will compare ¯$\bar{\mathcal{R}}(C_{P^{*}P})$ and ¯$\bar{\mathcal{R}}(\mathcal{G}(P^{*}P))$. Recall that ¯$\bar{\mathcal{R}}(\mathcal{G}(P^{*}P))$ is decided only by the graph topology of $\mathcal{G}(P^{*}P)$, and is independent of each element of $P^{*}P$. In contrast, ¯$\bar{\mathcal{R}}(C_{P^{*}P})$ is affected by each element of $C_{P^{*}P}$, which can vary to each other.
Therefore, we first consider the range of the elements of $C_{P^{*}P}$. When $(C_{P^{*}P})_{uv}\neq 0$, there is an index $w$ such that $P_{wu}$ and $P_{wv}$ are both positive, so it holds that $(C_{P^{*}P})_{uv}=n\sum_{w}\pi_{w}P_{wu}P_{wv}\geq n\pi_{\min}p_{\min}^{2}$. Moreover, the number of nonzero entries of $P_{wu}$ for any fixed $u$ does not exceed $\delta_{\mathrm{in}}+1$. Therefore, we obtain $(C_{P^{*}P})_{uv}\leq n\pi_{\max}(\delta_{\mathrm{in}}+1)p_{\max}^{2}$. Applying Lemma 3.14 to $(C_{P^{*}P})_{uv}$, we obtain
@@
Next, we evaluate ¯$\bar{\mathcal{R}}(\mathcal{G}(P^{*}P))$ using ¯$\bar{\mathcal{R}}(\mathcal{G}(P))$, both of which are determined by the positions of the nonzero elements in $P$. Because $(P^{*}P)_{uv}=\sum_{w}\pi_{u}^{-1}\pi_{w}P_{wu}P_{wv}$, $(P^{*}P)_{uv}$ is positive if and only if there exists an index $w$ which $P_{wu}$ and $P_{wv}$ are positive. From a graph theoretic perspective, this means that $\mathcal{G}(P^{*}P)$ has an edge $\{u,v\}$ if and only if there exists a node $w$ with edges $(w,u)$ and $(w,v)$ in $\mathcal{G}_{\mathrm{dir}}(P)$.
To compare the places of edges in $\mathcal{G}(P^{*}P)$ and $\mathcal{G}(P)$, we introduce the novel concepts of a back-and-forth path and its pivot node. A back-and-forth path between $u$ and $v$ is a path which consists of $\{u,w\},\{w,v\}\in\mathcal{G}(P)$ where $(w,u)$ and $(w,v)$ are edges in $\mathcal{G}_{\mathrm{dir}}(P)$. We refer to $w$ as the pivot node of the path (see Figure 2). When $\mathcal{G}_{\mathrm{dir}}(P)$ is symmetric, namely, $P_{ij}\neq 0$ if and only if $P_{ji}\neq 0$, back-and-forth paths reduce to paths of length $2$ in $\mathcal{G}(P)$. In such cases, including when $P$ is reversible, $\mathcal{G}(P^{*}P)$ has the same graph topology as $\mathcal{G}(P^{2})$. Generally, the graph $\mathcal{G}(P^{h})$ is called $h$-fuzz, and previous research has shown the relationship between ¯$\bar{\mathcal{R}}(\mathcal{G}(P))$ and ¯$\bar{\mathcal{R}}(\mathcal{G}(P^{h}))$ [2]. However, for nonreversible $P$, it is worthwhile to compare $\mathcal{G}(P)$ and $\mathcal{G}(P^{*}P)$.
Using the concept of a back-and-forth path, we can obtain an upper bound by a simple way. For each edge $(u,v)$ in $\mathcal{G}_{\mathrm{dir}}(P)$, we can choose a back-and-forth path between $u$ and $v$, because there are edges $(u,u)$ (a self loop) and $(u,v)$ in $\mathcal{G}_{\mathrm{dir}}(P)$. Therefore, any edge appearing in $\mathcal{G}(P)$ also appears in $\mathcal{G}(P^{*}P)$. By this property, we can obtain the following upper bound for ¯$\bar{\mathcal{R}}(\mathcal{G}(P^{*}P))$:
Lemma 3.15.
$\mathcal{R}_{uv}\left(\mathcal{G}(P)\right)\geq\mathcal{R}_{uv}\left(\mathcal{G}(P^{*}P)\right).$
Proof 3.16.
Because all edges in $\mathcal{G}(P)$ appear in $\mathcal{G}(P^{*}P)$, this lemma is a direct result from Rayleigh’s monotonicity law.
We can also obtain a lower bound for ¯$\bar{\mathcal{R}}(\mathcal{G}(P^{*}P))$. The following lemma is a variation of an argument for $h$-fuzz shown in Lemma 5.5.1. of [2].

### v2

@@ 2.4 The relationships between linear consensus algorithm and effective resistance
The proof is analogous to the argument of Section 3.1.2 in [21].
Let $\Phi(P):=\Phi_{n}(P)=n\varPi P$, where $n$ denotes the number of rows of $P$. Then we can see a clear relationship between the random walk with the transition matrix $P$ and some properties in the electrical network with conductance matrix $\Phi(P)$ such as the effective resistance, the voltage and the current (see [10]). In the analogy of random walk, a matrix called Green matrix [21] plays an important role. The Green matrix of consensus matrix $P$ is defined as:
@@
This matrix is also called the fundamental matrix [1]. For a consensus matrix $P$, this matrix is well defined (see Chapter 11.5 in [15]).
@@
Because $P^{t}-\bm{1}{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\bm{\pi}^{\mathsf{T}}}=(P-\bm{1}\bm{\pi})^{t}$ for $t\geq 1$, (7) can be reformulated as follows:
(8) $G(P)+\bm{1}{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\bm{\pi}^{\mathsf{T}}}=I+\sum_{t\geq 1}(P-\bm{1}\bm{\pi}^{\mathsf{T}})^{t}.$
@@
(9) $G(P)+\bm{1}{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\bm{\pi}^{\mathsf{T}}}=(I-P+\bm{1}\bm{\pi}^{\mathsf{T}})^{-1}.$
A fundamental matrix often refers not to $G(P)$, but to the right-hand side of (8) or (9) (e.g. [15, 19]).
The reason for introducing the Green matrix is that for reversible consensus matrix $P$, the effective resistance of $C:=\Phi(P)$ can be written as
$\mathcal{R}_{ab}(C)=\frac{1}{n}(\bm{e}_{a}-\bm{e}_{b})^{\mathsf{T}}G(P)\varPi^{-1}(\bm{e}_{a}-\bm{e}_{b}),$
which can be found in [21]. Such properties in this section are used to relate $J(P)$ and the effective resistance of the graph. In particular, $G(P)$ plays an important role in evaluating $J(P)$.
@@ 3 Performance analysis of linear consensus algorithm using effective resistance
In this section, we give upper and lower bounds for $J(P)$ using the concept of effective resistance. A bound for $J(P)$ has been known for reversible $P$ [21, 19], but not for nonreversible cases. To provide bounds for nonreversible $P$, we use a reversible matrix $P^{*}P$, where $P^{*}:=\varPi^{-1}P^{\mathsf{T}}\varPi$. Our result is a generalization of [21].
@@ 3.1 Properties of $P^{*}P$
Because $P$ is normal, $(I-P)(I-P^{\mathsf{T}})=(I-P^{\mathsf{T}})(I-P)$. Multiplying $\bm{1}$ from the right, we obtain $(I-P)(I-P^{\mathsf{T}})\bm{1}=\bm{0}$, which means $(I-P^{\mathsf{T}})\bm{1}\in\ker(I-P)$. Therefore, there exists some real value $a$ that satisfies $(I-P^{\mathsf{T}})\bm{1}=a\bm{1}$, because $\ker(I-P)=\operatorname{span}\{\bm{1}\}$. Then we multiply $\bm{1}^{\mathsf{T}}$ from the left and obtain $\bm{0}=an$. This means that $a=0$ and $(I-P^{\mathsf{T}})\bm{1}=\bm{0}$. Thus, $P$ is doubly-stochastic and the invariant measure $\bm{\pi}$ is $\frac{1}{n}\bm{1}$. Therefore, we obtain $P^{*}={\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}P^{\mathsf{T}}}$ and $P^{*}P=P^{\mathsf{T}}P=PP^{\mathsf{T}}=PP^{*}$.
@@
Inclusion properties among classes of consensus matrices are shown in Figure 1. Notice that there is also a matrix $P$ which is neither reversible, nor normal, but satisfies $P^{*}P=PP^{*}$, such as the following example (11). Thus, the condition $P^{*}P=PP^{*}$ is a weaker assumption than reversibility or normality.
@@
3.2 A bound for $J(P)$ and $J_{\mathrm{w}}(P)$ using $\bar{\mathcal{R}}(C)$
@@
(12) $\bar{\mathcal{R}}_{\mathrm{w}}(C):=\frac{1}{2}\bm{\pi}^{\mathsf{T}}\mathcal{R}(C)\bm{\pi}=\frac{1}{2}\sum_{(u,v)\in V\times V}\mathcal{R}_{uv}(C)\pi_{u}\pi_{v},$
@@
$\bar{\mathcal{R}}_{\mathrm{w}}(C)\leq\frac{1}{2}\pi_{\max}^{2}\sum_{(u,v)\in V\times V}\mathcal{R}_{uv}(C)=n^{2}\pi_{\max}^{2}\bar{\mathcal{R}}(C),$
@@
$\bar{\mathcal{R}}_{\mathrm{w}}(C)\geq\frac{1}{2}\pi_{\min}^{2}\sum_{(u,v)\in V\times V}\mathcal{R}_{uv}(C)=n^{2}\pi_{\min}^{2}\bar{\mathcal{R}}(C).$
@@
$\bar{\mathcal{R}}_{\mathrm{w}}(C)=\frac{1}{n}\operatorname{tr}G(P),$
@@
Using $\bar{\mathcal{R}}_{\mathrm{w}}(C)$, we prove Theorem 3.5.
@@
because ${\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}J_{\mathrm{w}}(P)}=\operatorname{tr}G_{\mathrm{w}}$ by definition. Since
@@
Lemmas 3.6 and 3.8 imply that
@@ 3.3 A bound for $J(P)$ using the graph structure of network
In this section, we establish bounds for $J(P)$ using the structure of associated undirected graph $\mathcal{G}(P)$, the maximum and minimum elements of $\bm{\pi}$, and the minimum element of $P$. Notice that we do not consider the edge weights of $\mathcal{G}(P)$ and focus only on its graph topology of $\mathcal{G}(P)$.
@@
Let $P$ be a consensus matrix with invariant measure $\bm{\pi}$, and let $\mathcal{G}(P)$ be the associated undirected graph with $P$. Then it holds that
@@
where $\pi_{\min}$ and $\pi_{\max}$ are respectively the minimum and maximum entries of $\bm{\pi}$, $p_{\min}$ is the minimum nonzero entry of $P$, and $\bar{\mathcal{R}}(\mathcal{G}(P))$ is the average effective resistance of a network, associated with $\mathcal{G}(P)$, such that each edge has unit conductance. Moreover, if $PP^{*}=P^{*}P$, then
$J(P)\geq\frac{\pi_{\min}^{3}n}{p_{\max}^{2}f(\delta_{\mathrm{in}})\pi_{\max}^{2}}\bar{\mathcal{R}}(\mathcal{G}(P)),\quad J_{\mathrm{w}}(P)\geq\frac{\pi_{\min}^{3}n^{2}}{p_{\max}^{2}f(\delta_{\mathrm{in}})\pi_{\max}}\bar{\mathcal{R}}(\mathcal{G}(P)).$
where $p_{\max}$ is the maximum nonzero entry of $P$, and $\delta_{\mathrm{in}}$ is the maximum indegree (excluding self loops) of the associated directed graph $\mathcal{G}_{\mathrm{dir}}(P)$, and $f(\delta_{\mathrm{in}}):=4\delta_{\mathrm{in}}^{2}+2\delta_{\mathrm{in}}-2$.
By Theorem 3.5, it is enough for the proof to show the following lemma, which is a generalization of Lemma 5.9 in [21].
@@
$\bar{\mathcal{R}}(C_{P^{*}P})\leq\frac{1}{n\pi_{\min}p_{\min}^{2}}\bar{\mathcal{R}}(\mathcal{G}(P)).$
@@
$\bar{\mathcal{R}}(C_{P^{*}P})\geq\frac{1}{n\pi_{\max}f(\delta_{\mathrm{in}})p_{\max}^{2}}\bar{\mathcal{R}}(\mathcal{G}(P)).$
Since the associated graph of the consensus matrix $P$ can be directed, we evaluate $\bar{\mathcal{R}}(C_{P^{*}P})$ using the maximum indegree $\delta_{\mathrm{in}}$ of the graph nodes of $\mathcal{G}_{\mathrm{dir}}(P)$, unlike previous research [21].
We employ the undirected graph $\mathcal{G}(P)$ instead of the directed graph $\mathcal{G}_{\mathrm{dir}}(P)$, because the following property is essential for proving Theorem 3.11.
@@
To prove Lemma 3.12, we first compare $\bar{\mathcal{R}}(C_{P^{*}P})$ and $\bar{\mathcal{R}}(\mathcal{G}(P^{*}P))$ using Lemma 3.13, and then compare $\bar{\mathcal{R}}(\mathcal{G}(P^{*}P))$ and $\bar{\mathcal{R}}(\mathcal{G}(P))$. Recall that $\bar{\mathcal{R}}(\mathcal{G}(P^{*}P))$ depends only on the graph topology of $\mathcal{G}(P^{*}P)$ and is independent of each element of $P^{*}P$. In contrast, $\bar{\mathcal{R}}(C_{P^{*}P})$ is affected by each element of $C_{P^{*}P}$, which can differ to each other.
Lemma 3.14.
Let $\mathcal{G}(P^{*}P)$ be the undirected graph associated with $P^{*}P$. Then, it holds that
@@
Proof 3.15.
We consider the range of the elements of $C_{P^{*}P}$. When $(C_{P^{*}P})_{uv}\neq 0$, there exists $w$ such that $P_{wu}$ and $P_{wv}$ are positive, and hence $(C_{P^{*}P})_{uv}=n\sum_{w}\pi_{w}P_{wu}P_{wv}\geq n\pi_{\min}p_{\min}^{2}$. Moreover, for any fixed $u$, the number of nonzero entries of $P_{wu}$ is at most $\delta_{\mathrm{in}}+1$. Therefore, we obtain $(C_{P^{*}P})_{uv}\leq n\pi_{\max}(\delta_{\mathrm{in}}+1)p_{\max}^{2}$. By applying Lemma 3.13 to $(C_{P^{*}P})_{uv}$, we obtain (18).
Next, we evaluate $\bar{\mathcal{R}}(\mathcal{G}(P^{*}P))$ using $\bar{\mathcal{R}}(\mathcal{G}(P))$, both of which are determined by the positions of the nonzero elements in $P$. Because $(P^{*}P)_{uv}=\sum_{w}\pi_{u}^{-1}\pi_{w}P_{wu}P_{wv}$, the entry $(P^{*}P)_{uv}$ is positive if and only if there exists $w$ such that both $P_{wu}$ and $P_{wv}$ are positive. From a graph-theoretic perspective, this means that $\mathcal{G}(P^{*}P)$ has an edge $\{u,v\}$ if and only if there exists a node $w$ with edges $(w,u)$ and $(w,v)$ in $\mathcal{G}_{\mathrm{dir}}(P)$.
To compare the places of edges in $\mathcal{G}(P^{*}P)$ and $\mathcal{G}(P)$, we introduce the novel concepts of a back-and-forth path and its pivot node. A back-and-forth path between $u$ and $v$ is a path which consists of $\{u,w\},\{w,v\}\in\mathcal{G}(P)$ where $(w,u)$ and $(w,v)$ are edges in $\mathcal{G}_{\mathrm{dir}}(P)$. We refer to $w$ as the pivot node of the path (see Figure 2). When $\mathcal{G}_{\mathrm{dir}}(P)$ is symmetric, namely, $P_{ij}\neq 0$ if and only if $P_{ji}\neq 0$, back-and-forth paths reduce to paths of length $2$ in $\mathcal{G}(P)$. In such cases, including when $P$ is reversible, $\mathcal{G}(P^{*}P)$ has the same graph topology as $\mathcal{G}(P^{2})$. In general, the graph $\mathcal{G}(P^{h})$ is called $h$-fuzz, and previous research has shown the relationship between $\bar{\mathcal{R}}(\mathcal{G}(P))$ and $\bar{\mathcal{R}}(\mathcal{G}(P^{h}))$ [2]. However, for nonreversible $P$, it is worthwhile to compare $\mathcal{G}(P)$ and $\mathcal{G}(P^{*}P)$. Notice that the concept of a pivot node is close to that of a common in-neighbor, but differs slightly. Here, we emphasize that both the pivot node and the back-and-forth are concepts on the undirected graph, although they are defined via conditions on the directed graph.
Using the concept of a back-and-forth path, we compare $\bar{\mathcal{R}}(\mathcal{G}(P^{*}P))$ and $\bar{\mathcal{R}}(\mathcal{G}(P))$ in the following lemma. Before stating the lemma, we recall that throughout the paper we assume that all diagonal entries of $P$ are positive. This ensures that every node in $\mathcal{G}_{\mathrm{dir}}(P)$ has a self-loop, which is used in the proof below.